# On the Utility of Accounting for Human Beliefs about AI Behavior in Human-AI Collaboration

### v1

@@ 5 Conclusion
Our work explores the impact of incorporating human behavior and beliefs into the design of collaborative AI, aiming to improve collaboration between humans and AI. By developing models that account for human beliefs regarding AI actions and integrating these models into AI design, we have observed improved performance in human-AI collaboration. Our approach, validated through simulations and human-subject experiments, demonstrates that AI agents designed with an understanding of human behavior and beliefs can be more effective in working with humans. This suggests a potential path forward for creating AI systems that are better aligned with human partners, making collaborative tasks more efficient and intuitive.
Generalization and limitations. We have demonstrated the effectiveness of our approach through human-subject experiments in commonly used grid world environments with varying levels of complexity. However, as with most human-subject studies, our findings are limited to the chosen environments. Moreover, our approach leverages historical data to construct models of human behavior and beliefs. This approach implicitly assumes that both human behavior and beliefs remain unchanged over time. Therefore, it is important to improve the development of human models and examine their generalizability to other environments.
Societal impacts. Our approach highlights the importance of incorporating not only human behavior but also human beliefs about AI behavior in the design of AI for human-AI collaboration. We believe this will become increasingly important as AI capabilities grow. However, while our research focuses on enhancing human-AI collaboration, there is a potential for the same methods to be used negatively, such as designing AI that intentionally sabotages human utility. One approach to mitigate potential negative impacts is to increase the transparency of AI models, enabling humans to develop appropriate reliance when working with AI agents.
Acknowledgements
@@
This work is supported in part by J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
@@ Appendix A Additional Experiments
Besides experiments in Section 4, we conduct additional experiments in a simple grid world environment with multiple goals to illustrate our study of humans’ beliefs about AI behavior.
A.1 Experiment Environment: Grid World with Multiple Goals
@@
Figure 6: Human-subject experiment interfaces. In Experiment 4, each participant is asked to control the player to move to the goal (red star). In Experiment 5, each participant is provided a trace of the behavior by another agent, and is asked to infer which goal the agent is trying to reach. In Experiment 6, each participant is playing with an AI agent in separate environments. The participants only receive bonus rewards by reaching the same goal (red star or green triangle) as the AI agent.
A.2 Experiment 4: Training and Evaluating Human Behavior Models in Single-Player MDP
@@
Table 3: The prediction accuracy for human behavior for different human models in Experiment 4.
@@
A.3 Experiment 5: Evaluating Human Belief Models
@@
Table 4: Performance comparison between Bayesian inference framework assuming standard model and human behavior model.
@@
Figure 7: Human evaluations of belief inference accuracy regrading AI goals.
@@
A.4 Experiment 6: Evaluating the Design of Collaborative AI Agents
@@
[…]
tableSimulation results of human-AI collaborative rewards in Experiment 6. Columns players are different AI agents, and row players are different simulated human models. Human Model AI Agent Self-play Behavior-AI Belief-AI Self-play AI 0.7828 0.4780 0.6164 Behavior 0.5926 0.7584 0.4552 Behavior&Belief 0.6919 0.7268 0.7813
Figure 8: Average collaborative reward in human-subject experiment 6.
@@ Appendix C Human Experiment Details
We include more details about human-subject experiments. The experiment interface is shown in Figure 9. We conduct four human-subject experiments, and recruited 1690 workers from Amazon Mechanical Turk in total. Table 5 lists the number of workers, the number of tasks per worker and the number of treatments in each experiment. Table 6 contains the demographic information of all the workers.
@@
Figure 9: Human-subject experiment interfaces.
Table 5: Number of workers in each human-subject experiment.
@@
Table 6: Demographic information of all the participants in our experiments.
@@
Age 20 to 29 775
30 to 39 642
40 to 49 146
50 or older 127
Gender Female 547
Male 1118
Other 25
Race / Ethnicity Caucasian 1511
Black or African-American 82
American Indian/Alaskan Native 32
Asian or Asian-American 28
Spanish/Hispanic 7
Other 30
Education High school degree 60
Some college credit, no degree 42
Associate’s degree 39
Bachelor’s degree 1217
Graduate’s degree 293
Other 39

### v2

@@ 5. Conclusion
Our work explores the impact of incorporating human beliefs about AI intentions into the design of collaborative AI, aiming to improve collaboration between humans and AI. By developing models that account for human beliefs regarding AI intentions and integrating these models into AI design, we have observed improved performance in human-AI collaboration. Our approach, validated through human-subject experiments, demonstrates that AI agents designed with an understanding of human behavior and beliefs over AI intention can be more effective in working with humans. This suggests a potential path forward for creating AI systems that are better aligned with human partners, making collaborative tasks more efficient and effective.
Generalization and limitations. Our approach has demonstrated effectiveness through human-subject experiments conducted in commonly used grid world environments of varying complexity. However, like most human-subject studies, our findings are limited to the chosen environments and conditions. To enhance the external validity of our approach, it is important to evaluate its performance across different domains and more realistic settings beyond grid worlds, such as dynamic, real-world applications involving physical robots or complex multi-agent systems.
Additionally, our approach leverages historical data to construct models of human behavior and beliefs, which implicitly assumes that human behavior and beliefs remain stable over time. This may not hold true in real-world scenarios where human behavior can change in response to prolonged exposure to AI or due to evolving task contexts. Future research could focus on developing adaptive models that can learn and update human behavior and belief patterns in real-time to address this limitation. Investigating the scalability of these models to larger, more diverse datasets and environments is also important and interesting for assessing the generalizability and robustness of our approach.
Societal impacts. Our work underscores the importance of integrating human behavior and beliefs about AI into the design of AI systems to enhance human-AI collaboration. As AI systems become increasingly autonomous and capable, it is critical that these systems are designed with an understanding of human behavior and beliefs. This alignment can help improve trust, reduce errors in collaborative settings, and foster more effective interactions.
However, there are also potential ethical considerations and potential risks due to the dual-use of our approaches. In particular, the same methods used to improve human-AI collaboration could be misused, such as designing AI systems that manipulate or deceive users by exploiting their expectations or beliefs. This could undermine trust and create systems that harm user welfare. To mitigate these risks, one potential approach is to develop and enforce ethical guidelines that prioritize transparency in AI models. Providing clear and understandable explanations for AI behavior can help users build appropriate levels of trust and reliance. Moreover, promoting responsible AI design principles that incorporate user consent, control, and oversight are also important in ensuring that the technology serves societal benefits rather than causing harm.
{acks}
@@
This work is supported in part by a J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
@@ Appendix A Additional Experiments
Besides experiments in Section 4, we also conduct additional experiments in a simple grid world environment with multiple goals to illustrate our study of humans’ beliefs about AI behavior. We also explore how to improve the model accuracy of behavioral cloning approach.
A.1. Experiment Environment: Grid World with Multiple Goals
@@
Figure 6. Human-subject experiment interfaces. In Experiment 4, each participant is asked to control the player to move to the goal (red star). In Experiment 5, each participant is provided a trace of the behavior by another agent, and is asked to infer which goal the agent is trying to reach. In Experiment 6, each participant is playing with an AI agent in separate environments. The participants only receive bonus rewards by reaching the same goal (red star or green triangle) as the AI agent.
A.2. Experiment 4: Evaluating Behavior Models in Single-Player MDP
@@
Table 3. The prediction accuracy for human behavior for different human models in Experiment 4.
@@
A.3. Experiment 5:
Evaluating Human Belief Models
@@ A.3. Experiment 5:
Table 4. Performance comparison between Bayesian inference framework assuming standard model and human behavior model.
@@
Figure 7. Human evaluations of belief inference accuracy regrading AI goals.
@@
A.4. Experiment 6:
Evaluating Collaborative AI Design
@@ A.4. Experiment 6:
Table 5. Simulation results of human-AI collaborative rewards in Experiment 6. Columns players are different AI agents, and row players are different simulated human models.
Human Model AI Agent
Self-play Behavior-AI Belief-AI
Self-play AI 0.7828 0.4780 0.6164
Behavior 0.5926 0.7584 0.4552
Behavior&Belief 0.6919 0.7268 0.7813
Figure 8. Average collaborative reward in human-subject experiment 6.
A.5. Extensive Results of Behavioral Cloning
We investigate techniques to enhance the accuracy of behavioral cloning models. Our experiment results reveal that human actions depend significantly on historical actions. We explicitly include the action history of both AI and human players as part of the model input to determine if this improves the prediction of human actions. The findings, presented in Figure 9, indicate that incorporating action history enhances the accuracy of our behavioral cloning model predictions. Additionally, our belief model consistently boosts prediction accuracy across nearly all scenarios.
(a) Behavioral cloning models.
(b) Behavioral cloning & belief models.
Figure 9. Evaluation of model accuracy on human-subject experiment data through incorporation of action history.
@@ Appendix B Implementation Details
Selection of population size.
For FCP and MEP algorithms used in Section 4, the population size will directly influence the model performance. We vary the number of agents in our simulations as shown in Figure 10. The model trained with 8 partners is chosen for use in further simulations and human-subject experiments for comparative analysis.
Figure 10. Average team reward for models trained across different population sizes using the FCP and MEP algorithms.
@@ Appendix C Human Experiment Details
We include more details about human-subject experiments. The experiment interface is shown in Figure 11. We conduct four human-subject experiments, and recruited 1990 workers from Amazon Mechanical Turk in total. Table 6 lists the number of workers, the number of tasks per worker and the number of treatments in each experiment. Table 7 contains the demographic information of all the workers.
@@
Figure 11. Human-subject experiment interfaces.
Table 6. Number of workers in each human-subject experiment.
@@
Table 7. Demographic information of all the participants in our experiments.
@@
Age 20 to 29 714
30 to 39 974
40 to 49 176
50 or older 72
Other 54
Gender Female 639
Male 1319
Other 32
Race / Ethnicity Caucasian 1772
Black or African-American 105
American Indian/Alaskan Native 36
Asian or Asian-American 29
Spanish/Hispanic 11
Other 37
Education High school degree 66
Some college credit, no degree 22
Associate’s degree 36
Bachelor’s degree 1457
Graduate’s degree 353
Other 56